# SDC-Net: A Domain Adaptation Framework with Semantic-Dynamic Consistency for Cross-Subject EEG Emotion Recognition

### v1

@@ Abstract
Electroencephalography(EEG) based emotion recognition holds great promise for affective brain-computer interfaces (aBCIs), yet practical deployment remains challenging due to substantial inter-subject variability and the lack of labeled data in target domains. To overcome these limitations, we present a novel unsupervised Semantic-Dynamic Consistency domain adaptation network for fully label-free cross-subject EEG emotion recognition. First, we introduce a Same-Subject Same-Trial Mixup strategy that generates augmented samples via intra-trial interpolation, enhancing data diversity while explicitly preserving individual identity to mitigate label ambiguity. Second, we construct a dynamic distribution alignment module in reproducing kernel Hilbert space (RKHS), jointly aligning marginal and conditional distributions through multi-objective kernel mean embedding, and leveraging a confidence-aware pseudo-labeling strategy to ensure stable adaptation. Third, we propose a dual-domain similarity consistency learning mechanism that enforces cross-domain structural constraints based on latent pairwise similarities, enabling semantic boundary learning without relying on temporal synchronization or label priors. To validate the effectiveness and robustness of the proposed SDC-Net, extensive experiments are conducted on three widely used EEG benchmark datasets: SEED, SEED-IV, and Faced. Comparative results against existing unsupervised domain adaptation methods demonstrate that SDC-Net achieves state-of-the-art performance in emotion recognition under both cross-subject and cross-session conditions. This advancement significantly improves the accuracy and generalization capability of emotion decoding, and lays a solid foundation for real-world applications of personalized affective brain-computer interfaces (aBCIs). The source code will be released at https://github.com/XuanSuTrum/SDC-Net.
Index Terms:
affective brain-computer interfaces, emotion recognition, electroencephalogram,domain adaptation, individual variability, transfer learning, Hilbert Space Embeddings
@@ I Introduction
Emotion play a pivotal role in human experience, reflecting the complex and dynamic interplay between physiological states and neural activity. Understanding the mechanisms underlying emotion generation, perception, and regulation is critical for advancing mental health interventions and enhancing human–computer interaction systems [1]. Against this backdrop, electroencephalography (EEG)-based emotion recognition has attracted considerable attention in affective computing due to its non-invasiveness and high temporal resolution [2, 3]. EEG signals capture rapid neural dynamics associated with emotional states, offering a valuable window into the underlying mechanisms of emotional processing. Furthermore, decoding emotion-related EEG patterns lays the theoretical foundation for developing emotion-aware brain–computer interfaces (BCIs) and adaptive human-centric systems.
Despite its promising potential, EEG-based emotion recognition still faces two primary challenges. First, EEG signals exhibit significant inter-subject variability and intra-subject non-stationarity, severely limiting the generalizability of recognition models across individuals. Second, acquiring high-quality emotion labels requires complex experimental protocols and substantial manual effort, which constrains the scalability of labeled datasets. These issues hinder the real-world deployment of EEG-based emotion recognition systems. Therefore, it is crucial to develop label-efficient and subject-robust recognition frameworks.
Transfer learning offers a compelling solution to these challenges by utilizing knowledge from a labeled source domain to boost performance in an unlabeled target domain[4]. In EEG-based emotion recognition, the source domain typically consists of labeled EEG recordings from multiple subjects, while the target domain comprises unlabeled EEG signals[5]. Transfer learning, particularly domain adaptation (DA), holds great potential to mitigate inter-subject variability by aligning distributions between domains and transferring learned representations from $D_{s}$ to $D_{t}$[6].
@@
Existing state-of-the-art DA methods for EEG-based emotion recognition can be broadly categorized into three paradigms. The first line of work enforces consistency constraints at the output or feature level to improve cross-domain generalization. Representative approaches include multi-source feature extraction [7, 8], adversarial feature alignment (e.g., DANN [9]), and pseudo-label propagation strategies [10]. While effective in aligning distributions, these approaches often ignore the semantic structure between instances and are prone to noisy pseudo-labels, particularly under unsupervised target settings. To further improve the discriminative capacity of the learned features, a second category of methods introduces structural modeling techniques, such as pairwise or triplet loss constraints [11, 12, 13, 14] and prototype-based representations. These methods typically guide target samples toward pre-learned emotion category prototypes in the source domain, treating them as ”semantic anchors.” However, due to distributional shifts, these anchors may become misaligned in the target domain, leading to suboptimal guidance and reduced transfer performance. More recently, contrastive learning has gained popularity for learning subject-invariant and emotion-discriminative representations by maximizing agreement between samples from the same emotional state and minimizing it across different states [15, 16, 17, 18]. Nevertheless, most contrastive frameworks rely on temporally synchronized experimental protocols to define positive and negative pairs, which implicitly leverage label information and limit their applicability in real-world unsupervised scenarios.
@@
To address the aforementioned challenges, we propose a novel domain adaptation framework for cross-subject EEG emotion recognition, which systematically tackles individual variability, semantic structure modeling, and distribution alignment. The main contributions of this work are as follows:
(1) Same-Subject Same-Trial Mixup (SS-Mix): Inspired by the Mixup augmentation strategy, we design an intra-subject, intra-trial sample mixing mechanism to enhance data diversity while preserving individual-specific traits. This strategy effectively mitigates the ambiguity between subject-specific features and emotion labels.
(2) Dynamic Distribution Alignment in RKHS: We construct a unified kernel mean embedding framework that jointly aligns marginal and conditional distributions in a shared RKHS space. This objective is formulated as a multi-objective optimization problem to adaptively balance subject-level (global) and emotion-level (semantic) alignment. A dynamic confidence-based mechanism is further introduced to progressively filter high-confidence pseudo-labeled samples, improving conditional alignment reliability and transfer robustness.
(3) Dual-Domain Similarity Consistency Learning(DSCL): We introduce a structure-aware constraint that enforces pairwise similarity consistency across both source and target domains, enabling the model to capture fine-grained semantic boundaries without relying on temporal synchronization. This enhances generalization to complex and unlabeled emotional scenarios.
Notations and descriptions used in this paper is shown in Table I.
@@
$x_{s}^{l}/x_{t}^{u}$ Source///Target domain samples
@@
$RKHS$ Reproducing Kernel Hilbert Space
$SGD$ Stochastic Gradient Descent
$ReLU$ Rectified Linear unit activation function
Figure 1: Traditional machine learning models are trained in isolation on separate datasets, without benefiting from shared knowledge. Transfer learning bridges this gap by reusing the knowledge learned by a source model, trained on a source dataset, and applying it to a target dataset, thereby improving efficiency. Within transfer learning, domain adaptation specifically addresses domain shift, a scenario where the data distributions of $D_{s}$ and $D_{t}$ are misaligned. By aligning these distributional differences, domain adaptation methods enable models to perform robustly on $D_{t}$, even with minimal labeled data.
II Releated Work
II-A EEG-Based Emotion Recognition and Deep Learning Approaches
Traditional EEG-based emotion recognition primarily relies on classical machine learning methods, which typically involve two separate stages: manual feature extraction followed by classification. These approaches depend heavily on handcrafted features, thus requiring researchers to possess strong domain expertise to identify emotion-relevant neural patterns. However, when applied to large-scale and complex EEG datasets, such methods often suffer from limited scalability and poor generalization performance[22, 23, 24].
With the advancement of deep learning, significant progress has been made in EEG-based emotion recognition. In particular, architectures such as Convolutional Neural Networks (CNNs) and Bidirectional Long Short-Term Memory networks (BiLSTMs) have demonstrated strong capabilities in automatically extracting discriminative features from raw or minimally preprocessed EEG signals[25]. For instance, certain methods construct three-dimensional EEG representations by integrating spatial and spectral information across channels, enabling CNN-based models to achieve arousal and valence classification accuracies of 89.67% and 90.93%, respectively, on benchmark datasets [26]. Furthermore, hybrid architectures combining CNNs and BiLSTMs have been explored, which simultaneously capture the spatial structure and temporal dependencies in EEG signals, thereby enhancing emotion recognition performance [24].
Despite the promising performance of deep learning methods in within-subject emotion recognition tasks, their generalization to cross-subject and cross-session scenarios remains a critical challenge due to individual differences and temporal variability in EEG signals. This limitation significantly restricts the applicability of these models in real-world scenarios [27, 28].
II-B Data Augmentation for EEG Emotion Recognition
Due to the scarcity and high acquisition cost of EEG signals, especially in emotion recognition tasks, recent work has turned to data augmentation (DA) techniques to improve model generalization. A representative line of research leverages generative adversarial networks (GANs) to synthesize high-quality EEG-like signals. For instance, GANSER[19] introduces a self-supervised GAN framework to generate diverse EEG samples by masking and regenerating partial inputs, thereby enriching the training distribution. Similarly, SA-cWGAN[20] incorporates self-attention to capture long-range dependencies in EEG signals and employs Wasserstein loss with gradient penalty to enhance sample fidelity, demonstrating notable gains on SEED and DEAP datasets. Du et al.[21] further propose L-C-WGAN-GP, which integrates LSTM-based generation with CNN-based discrimination, producing EEG signals that statistically resemble real distributions across frequency and time domains.
However, most GAN-based augmentation methods focus on enhancing sample quantity rather than preserving semantic consistency across individuals. These methods often neglect the intrinsic variability in emotional EEG responses across subjects, which can unintentionally exacerbate the inter-subject discrepancy. Moreover, direct synthetic sampling may distort class boundaries or introduce ambiguity between subject-specific features and emotional semantics.
II-C Domain Adaptation Methods for Cross-Subject EEG Emotion Recognition
Electroencephalography (EEG) has been widely adopted in affective computing due to its high temporal resolution and direct access to cognitive and emotional states. However, the presence of significant inter-subject variability and inherent nonstationarity limits the generalization capability of emotion recognition models across individuals. To address this, DA techniques have been actively explored as effective strategies to transfer knowledge from labeled source domains to unlabeled or sparsely labeled target domains. A conceptual overview illustrating the differences among traditional learning, transfer learning, and domain adaptation in cross-domain EEG settings is presented in  Fig. 1. This figure highlights the core challenges of distribution mismatch and label scarcity that domain adaptation aims to address. Building upon this framework, existing DA methods for EEG-based emotion recognition can be broadly categorized into three paradigms.
II-C1 Feature Alignment and Adversarial Methods
A major line of DA research focuses on aligning feature distributions between domains to learn domain-invariant representations. Early approaches employed statistical alignment techniques or adversarial objectives to reduce marginal and conditional distributional shifts. For example, Chen et al.[7], and Li et al.[8] leveraged multi-source domain adaptation frameworks to extract both shared and domain-specific features. Adversarial strategies, such as DANN[9] and PR-PL[13], introduced domain discriminators trained jointly with the feature extractor to enforce domain confusion, promoting feature-level invariance. Additionally, pseudo-labeling and consistency regularization techniques have been adopted to refine target predictions in an unsupervised manner[10]. While these methods are effective in reducing global distribution discrepancies, they often lack mechanisms to model semantic relationships between instances, making them vulnerable to label noise and impairing their ability to delineate fine-grained emotional states.
II-C2 Similarity-based and Instance-level Methods
To overcome the limitations of purely distribution-based alignment, another line of research integrates structural modeling techniques to enhance the semantic discriminability of EEG features across domains. These methods typically incorporate metric-based constraints, such as pairwise or triplet losses, to enforce intra-class compactness and inter-class separability in the latent space [11, 12]. In addition, prototype-based methods utilize category centroids from the source domain as semantic anchors, encouraging target samples to align with these prototypical representations [13, 14]. While effective under moderate domain shift, this alignment often becomes unreliable in fully unsupervised settings due to the misalignment of prototypes caused by distributional discrepancies. Moreover, these approaches typically depend on accurate source label structures and can be sensitive to noisy or incorrect pseudo-labels when applied to unlabeled target domains.
Prototype-based methods treat the centroid of each emotion category in the source domain as a ”semantic anchor” and encourage target samples to align with their nearest prototypes in the latent space. However, under domain shifts, these prototypes are prone to positional drift in the target domain, which undermines their effectiveness as guidance signals. Moreover, such methods rely heavily on accurate source label structures and are susceptible to noisy pseudo-labels when applied in low-label or fully unsupervised settings.
II-C3 Contrastive Learning Methods
Contrastive learning has emerged as a powerful paradigm for learning discriminative and subject-invariant EEG representations in emotion recognition tasks without relying on explicit supervision. These methods typically construct positive and negative sample pairs based on the assumption that EEG signals elicited under the same emotional stimuli should be mapped closer in the feature space, while signals from different stimuli should remain apart. Shen et al.[15] proposed CLISA, which aligns inter-subject EEG representations using a contrastive objective, treating samples from the same video as positive pairs. Similarly, Dai et al.[16] incorporated spatial constraints across brain areas to generate region-aware contrastive pairs, improving intra-region consistency. Hu et al.[17] extended this idea by computing cross-subject correlations in temporal, spatial, and frequency domains, aligning representations using a correlation-based contrastive loss. Wang et al.[18] introduced GC-STCL, which employed Granger causality to enhance spatial-temporal structure learning in contrastive settings.
Despite achieving promising results, most of these methods define contrastive pairs based on temporally synchronized protocols—such as subjects watching the same video simultaneously—which implicitly encode label-related information. This contradicts the assumptions of fully unsupervised learning and limits generalizability in real-world, asynchronous, or spontaneous emotional environments. Moreover, their reliance on temporal alignment or external synchronization protocols constrains their applicability in more naturalistic or deployment settings.
III METHODOLOGY
As illustrated in Fig. 2, the proposed SDC-Net framework is designed to address the challenges of cross-subject EEG emotion recognition under fully label-free conditions. The overall pipeline begins with EEG data augmentation and feature extraction (Fig. 2a), where we introduce a SS-Mix strategy to improve the diversity and robustness of input samples. This augmentation is performed independently on both source and target domains, ensuring that subject-specific traits are preserved while mitigating overfitting caused by limited data. The augmented EEG signals are then passed through a shared feature extractor to obtain high-level representations. To reduce the domain discrepancy between source and target subjects, we propose a dynamic distribution alignment mechanism in RKHS (Fig. 2b). Specifically, we jointly align the marginal and conditional distributions using MMD and CMMD, formulated as a multi-objective optimization problem. A dynamic thresholding strategy is further applied to progressively select high-confidence pseudo-labeled target samples, improving the reliability of conditional alignment and enhancing transfer robustness. Finally, to capture fine-grained semantic relationships beyond global alignment, we introduce a DSCL strategy (Fig. 2c). This component enforces structure-aware similarity consistency between paired samples across domains by aligning pairwise distances in both feature and label spaces. Importantly, this strategy does not rely on temporal synchronization, allowing the model to generalize effectively to complex and unlabeled emotional scenarios. Together, these components form a unified and fully unsupervised domain adaptation framework tailored for robust EEG-based emotion recognition across subjects.
@@ III METHODOLOGY
III-A Data augmentation and Feature Extractor
III-A1 EEG data augmentation Based SS-Mix
The Mixup method performs data augmentation by creating convex combinations of pairs of samples, which extends the distribution space of the training data to a certain extent and enhances the model’s generalization capability. This simple yet effective method[29] has been proven to possess unique advantages in improving model robustness and reducing overfitting.
@@ III-A1 EEG data augmentation Based SS-Mix
where $\left(x_{i},y_{i}\right)$ and $\left(x_{j},y_{j}\right)$ are two examples drawn at random from the training data, and $\lambda$ is a random variable drawn from a Beta distribution $\left(\alpha,\alpha\right)$. $\alpha$ is a hyperparameter that controls the strength of the interpolation.
Based on the aforementioned approach, we further refined the data augmentation strategy by individually extracting and augmenting data from from the same subject and the same trial, rather than augmenting all data together. This method better preserves individual differences across trials within each subject and ensures that the augmented samples maintain consistency and physiological relevance.
Specifically, we segmented each subject’s dataset into different trials and applied the Mixup method separately to each trial. Within each trial, sample pairs were augmented as follows:
@@
Here, $x_{trial}^{i}$ and $x_{trial}^{j}$ are two randomly selected samples within the same trial, $y_{trial}^{i}$ and $y_{trial}^{j}$ are their corresponding labels. Unlike global data augmentation, this method ensures the consistency of data within trials while further enhancing the preservation of individual differences during the augmentation process. This approach not only expands the distributional space of the data but also minimizes potential physiological and psychological differences between subjects that could interfere with model training. Consequently, it enhances the model’s generalization capability and robustness in cross-subject emotion recognition tasks.
III-A2 Feature Extractor
In the domain of EEG signal analysis, a rigorous feature extraction process was implemented to distill meaningful information from the raw EEG data. The procedure began with the careful selection of specific time segments from the EEG signal, aiming to capture temporal information relevant to the analysis. Following this, the short-time Fourier transform (STFT) was employed to decompose the EEG signal into its constituent frequency bands: delta ($\delta$), theta ($\theta$), alpha ($\alpha$), beta ($\beta$), and gamma ($\gamma$) bands. This decomposition allowed for the isolation and examination of different frequency components, each of which is associated with distinct neurological and cognitive processes.
Within each frequency band, the probability density function (PDF) of the signal amplitude was calculated. The PDF provides a statistical representation of the distribution of signal amplitudes within each frequency band, offering insights into the underlying characteristics of the EEG signal.
The central aspect of this feature extraction process is the computation of differential entropy, a measure that quantifies the uncertainty and complexity of the EEG signal. Differential entropy is a continuous analog of the discrete Shannon entropy and is defined as:
[…]
In Eq.(5), $H(X)$ represents the differential entropy of the random variable $x$, which in this context corresponds to the amplitude of the EEG signal. The function $p(x)$ denotes the probability density function of the amplitude distribution within a given frequency band. The differential entropy $H(X)$ thus provides a scalar measure of the unpredictability or information content inherent in the EEG signal within that specific frequency band.
This feature extraction process was systematically applied to each frequency band, computing differential entropy values for each individual band. As a result, each EEG sample is represented by a differential entropy feature vector of dimension $N_{cf}=c\times f$, where $c$ denotes the number of EEG channels and $f=5$ corresponds to the number of frequency bands. This representation effectively captures the spectral-spatial characteristics of the EEG signals, providing a robust and informative foundation for further analysis and interpretation.
III-B Dynamic Distribution Alignment in RKHS

### v2

@@ Abstract
Emotion recognition based on electroencephalography (EEG) holds significant promise for affective brain-computer interfaces (aBCIs). However, its practical deployment faces challenges due to the variability within inter-subject and the scarcity of labeled data in target domains. To overcome these limitations, we propose SDC-Net, a novel Semantic-Dynamic Consistency domain adaptation network for fully label-free cross-subject EEG emotion recognition. First, we introduce a Same-Subject Same-Trial Mixup strategy that generates augmented samples through intra-trial interpolation, enhancing data diversity while explicitly preserving individual identity to mitigate label ambiguity. Second, we construct a dynamic distribution alignment module within the Reproducing Kernel Hilbert Space (RKHS), jointly aligning marginal and conditional distributions through multi-objective kernel mean embedding, and leveraging a confidence-aware pseudo-labeling strategy to ensure stable adaptation. Third, we propose a dual-domain similarity consistency learning mechanism that enforces cross-domain structural constraints based on latent pairwise similarities, facilitating semantic boundary learning without reliance on temporal synchronization or label priors. To validate the effectiveness and robustness of the proposed SDC-Net, extensive experiments are conducted on three widely used EEG benchmark datasets: SEED, SEED-IV, and FACED. Comparative results against existing unsupervised domain adaptation methods demonstrate that SDC-Net achieves state-of-the-art performance in emotion recognition under both cross-subject and cross-session conditions. This advancement significantly improves the accuracy and generalization capability of emotion decoding, laying a solid foundation for real-world applications of personalized aBCIs. The source code is available at: https://github.com/XuanSuTrum/SDC-Net.
Preprint Notice: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Emotion is fundamental to human experience, reflecting the interplay between physiological states and neural activity. With advances in affective computing, emotional states are increasingly quantified as measurable variables, driving emotion-aware systems in medical, industrial, and consumer domains[1, 2]. Intelligent human–machine interaction further underscores the need for nuanced emotional perception in applications such as rehabilitation robotics, health monitoring, and affective companionship[3, 4]. Future systems must move beyond external behaviors toward internal states to enable natural and personalized interaction.
Emotion recognition draws on behavioral and physiological cues, including facial expressions, gestures, speech, and biosignals[2]. Physiological signals—such as ECG, EMG, EOG, respiration, and EEG—are particularly robust due to their objectivity[5]. Among them, EEG is favored for its non-invasive nature and high temporal resolution, enabling real-time emotion monitoring in brain–computer interfaces and adaptive systems. Despite progress, two challenges persist: (1) large inter-subject variability and intra-subject non-stationarity hinder generalization; (2) acquiring reliable emotion labels is labor-intensive, limiting dataset scalability. These issues restrict real-world deployment, highlighting the need for label-efficient and subject-robust methods. Transfer learning addresses these challenges by leveraging labeled data from a source domain to improve performance in an unlabeled target domain[6, 7]. Domain adaptation (DA) in particular reduces inter-subject variability by aligning distributions and transferring representations between domains[8].
@@
Existing state-of-the-art DA methods for EEG-based emotion recognition can be broadly categorized into three paradigms. The first line of work enforces consistency constraints at the output or feature level to improve cross-domain generalization. Representative approaches include multi-source feature extraction [9, 10], adversarial feature alignment (e.g., DANN [11]), and pseudo-label propagation strategies [12]. While effective in aligning distributions, these approaches often overlook inter-instance semantic structures and are vulnerable to noisy pseudo-labels, especially in unsupervised target settings. To address discriminative limitations, a second category of methods introduces structural modeling techniques, such as pairwise or triplet loss constraints [13, 14, 15, 16] and prototype-based representations. These methods typically guide target samples toward pre-learned emotion category prototypes in the source domain, treating them as ”semantic anchors.” However, distributional shifts can misalign these anchors in the target domain, yielding suboptimal guidance and diminished transfer performance. More recently, contrastive learning has gained popularity for learning subject-invariant and emotion-discriminative representations by maximizing agreement between samples from the same emotional state and minimizing it across different states [17, 18, 19, 20]. Nevertheless, most contrastive frameworks rely on temporally synchronized experimental protocols to define positive and negative pairs, which implicitly leverage label information and limit their applicability in real-world unsupervised scenarios. A more comprehensive review of related studies, including deep learning, data augmentation, and domain adaptation methods for EEG-based emotion recognition, is provided in Supplementary Information (Section S5).
@@
To overcome these limitations, we propose a novel domain adaptation framework for cross-subject EEG emotion recognition, which systematically tackles individual variability, semantic structure modeling, and distribution alignment. The main contributions of this work are as follows:
(1) Same-Subject Same-Trial Mixup (SS-Mix): Inspired by the Mixup augmentation strategy[24], we design an intra-subject, intra-trial sample mixing mechanism to enhance data diversity while preserving individual-specific traits. This strategy effectively mitigates the ambiguity between subject-specific features and emotion labels.
(2) Dynamic Distribution Alignment (DDA) in RKHS: We construct a unified kernel mean embedding framework that jointly aligns marginal and conditional distributions in a shared RKHS space. This objective is formulated as a multi-objective optimization problem to adaptively balance subject-level (global) and emotion-level (semantic) alignment. A dynamic confidence-based mechanism is further introduced to progressively filter high-confidence pseudo-labeled samples, improving conditional alignment reliability and transfer robustness.
(3) Dual-Domain Similarity Consistency Learning (DSCL): We propose a structure-aware constraint that enforces pairwise similarity consistency across both source and target domains. This approach enables the model to effectively capture fine-grained semantic boundaries without relying on temporal synchronization. Consequently, it enhances generalization capabilities in complex and unlabeled emotional scenarios. Notations and descriptions used in this paper is shown in TableI.
@@
$x_{s}^{l}/x_{t}^{u}$ Source//Target domain samples
@@
RKHSRKHS Reproducing Kernel Hilbert Space
SGDSGD Stochastic Gradient Descent
ReLUReLU Rectified Linear unit activation function
II METHODOLOGY
@@
II-A Data augmentation and Feature Extractor
II-A1 EEG data augmentation Based SS-Mix
The Mixup method performs data augmentation by creating convex combinations of pairs of samples, which extends the distribution space of the training data to a certain extent and enhances the model’s generalization capability. This simple yet effective method has been proven to possess unique advantages in improving model robustness and reducing overfitting.
@@ II-A1 EEG data augmentation Based SS-Mix
where $\left(x_{i},y_{i}\right)$ and $\left(x_{j},y_{j}\right)$ are two examples selected randomly from the training data.
Based on the aforementioned approach, we further refined the data augmentation strategy by individually extracting and augmenting data from the same subject and the same trial, rather than augmenting all data together. This method better preserves individual differences across trials within each subject and ensures that the augmented samples maintain consistency and physiological relevance.
Specifically, we proposed a trial-wise augmentation strategy, termed SS-Mix, where each subject’s dataset was segmented into individual trials, and the Mixup method was applied separately within each trial. In this setting, sample pairs were augmented as follows:
@@
Here, $x_{trial}^{i}$ and $x_{trial}^{j}$ are two randomly selected samples within the same trial, $y_{trial}^{i}$ and $y_{trial}^{j}$ are their corresponding labels, where $\omega$ is a random mixing coefficient drawn from a Beta distribution over the interval $(0,1)$. Unlike global data augmentation, this method ensures the consistency of data within trials while further enhancing the preservation of individual differences during the augmentation process. This approach not only expands the distributional space of the data but also minimizes potential physiological and psychological differences between subjects that could interfere with model training. Consequently, it enhances the model’s generalization capability and robustness in cross-subject emotion recognition tasks.
II-A2 Feature Extractor
To extract informative representations from raw EEG signals, we first applied short-time Fourier transform (STFT) to decompose the data into five canonical frequency bands: delta ($\delta$), theta ($\theta$), alpha ($\alpha$), beta ($\beta$), and gamma ($\gamma$). For each band, we estimated the probability density function (PDF) of the amplitude distribution and computed its differential entropy (DE):
$H(x)=-\int p(x)\log p(x)dx,$ (5)
where $p(x)$ denotes the PDF. DE provides a compact measure of the information content within each band. Concatenating the DE features across all channels and frequency bands yields a spectral–spatial feature vector of dimension $N_{cf}=c\times f$ ($c$: channels, $f=5$: frequency bands), which serves as the input to the network.
II-B Dynamic Distribution Alignment in RKHS